# The Dial-a-Ride Problem with Limited Pickups per Trip

### v2

@@ Appendix B Full computational results for DARP-LPT
Table 1: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Basic instances, $L=4$)
@@
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 73 0.2 0.3 1.0 295.8 295.8 0.00% 64 0.1 0.0 0.2 295.8 295.8 0.00% 64 0.1 0.0 0.2 295.8 295.8 0.00% a3-30 93 0.3 17.5 18.9 477.0 477.0 0.00% 89 0.1 4.5 4.7 477.0 477.0 0.00% 89 0.2 4.4 4.7 477.0 477.0 0.00% a3-36 121 0.1 534.4 535.0 549.8 549.8 0.00% 115 0.2 36.0 36.4 549.8 549.8 0.00% 120 0.2 200.2 200.6 549.8 549.8 0.00% a4-48 318 0.3 1798.6 1800.0 669.4 662.9 0.97% 342 0.7 1798.9 1800.0 669.4 664.4 0.74% 305 0.7 1798.3 1800.0 669.4 665.5 0.58% a6-60 601 0.4 1798.0 1800.0 825.0 819.0 0.72% 817 4.5 1796.3 1800.0 825.0 820.3 0.57% 589 3.6 1794.6 1800.0 825.0 820.6 0.54% a7-70 1036 0.7 1793.6 1800.0 926.7 912.4 1.54% 2114 10.8 1786.0 1800.0 927.6 912.8 1.60% 944 10.7 1786.6 1800.0 927.6 913.1 1.56% a8-80 1088 0.7 1790.5 1800.0 1000.9 973.1 2.77% 3020 24.2 1770.1 1800.0 991.3 973.1 1.83% 1071 25.3 1767.5 1800.0 992.8 974.0 1.89% Avg-A 475.7 0.4 1107.8 0.86% 937.3 5.8 1034.5 0.68% 454.6 5.8 1057.9 0.65% b3-36 169 0.1 22.7 23.1 557.9 557.9 0.00% 210 0.3 13.5 14.0 557.9 557.9 0.00% 154 0.3 16.2 16.8 557.9 557.9 0.00% b4-40 311 0.2 1799.3 1800.0 627.4 623.9 0.56% 228 1.5 1696.5 1698.5 627.4 627.4 0.00% 216 1.7 1797.9 1800.0 627.4 626.2 0.20% b5-50 518 0.3 1798.7 1800.0 751.9 745.3 0.87% 1868 5.9 1791.6 1800.0 751.9 744.2 1.03% 492 7.1 1793.4 1800.0 751.9 745.7 0.82% b6-60 940 0.6 1795.9 1800.0 877.0 868.0 1.03% 3769 14.7 1782.2 1800.0 877.0 867.1 1.13% 806 14.7 1783.4 1800.0 877.0 867.8 1.05% b7-70 960 0.7 1798.2 1800.0 896.4 892.9 0.39% 6171 28.5 1766.2 1800.0 896.4 894.6 0.21% 891 31.7 1765.5 1800.0 896.4 894.8 0.19% Avg-B 579.6 0.4 1444.6 0.57% 2449.2 10.2 1422.5 0.47% 511.8 11.1 1443.4 0.45% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
@@
Table 2: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Basic instances, $L=6$)
@@
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 96 0.1 0.3 0.4 288.5 288.5 0.00% 81 0.1 0.5 0.6 288.5 288.5 0.00% 81 0.1 0.5 0.7 288.5 288.5 0.00% a3-30 106 0.1 1799.8 1800.0 491.4 486.6 0.97% 101 0.1 1799.7 1800.0 491.4 485.9 1.12% 101 0.1 1799.7 1800.0 491.4 486.7 0.96% a3-36 161 0.1 1799.5 1800.0 541.9 532.2 1.79% 159 0.2 1798.7 1800.0 541.9 535.4 1.20% 155 0.2 1799.7 1800.0 541.9 536.4 1.01% a4-48 420 0.2 1798.2 1800.0 697.8 680.2 2.52% 431 0.6 1798.0 1800.0 697.8 680.9 2.42% 407 0.7 1798.6 1800.0 697.9 679.9 2.58% a6-60 859 0.6 1793.8 1800.0 846.1 828.9 2.04% 1026 2.3 1794.7 1800.0 847.5 830.6 1.99% 830 2.2 1793.9 1800.0 846.1 826.6 2.31% a7-70 1502 1.2 1792.9 1800.0 971.5 939.8 3.26% 2055 5.0 1789.1 1800.0 968.3 941.4 2.78% 1285 6.8 1786.7 1800.0 968.9 940.5 2.93% a8-80 1483 1.1 1788.7 1800.0 1073.8 1004.4 6.46% 2274 7.0 1782.7 1800.0 1065.2 1006.0 5.56% 1430 10.7 1771.7 1800.0 1060.3 1003.0 5.41% Avg-A 661.0 0.5 1542.9 2.44% 875.3 2.2 1542.9 2.15% 612.7 3.0 1543.0 2.17% b3-36 460 0.3 1799.4 1800.0 556.5 540.9 2.80% 567 0.6 1799.1 1800.0 556.5 542.9 2.45% 435 0.7 1797.7 1800.0 556.5 544.0 2.25% b4-40 910 0.6 1798.1 1800.0 582.8 571.0 2.02% 906 2.1 1797.4 1800.0 582.8 573.4 1.61% 614 1.8 1797.2 1800.0 582.8 573.8 1.55% b5-50 1686 1.4 1792.0 1800.0 755.1 730.2 3.30% 3196 12.4 1785.3 1800.0 754.7 732.3 2.97% 1348 7.7 1789.4 1800.0 754.7 731.5 3.07% b6-60 3036 2.9 1790.3 1800.0 873.8 837.0 4.21% 3890 9.6 1781.1 1800.0 868.9 841.7 3.13% 1929 10.5 1782.0 1800.0 874.6 842.0 3.73% b7-70 3478 3.4 1753.1 1800.0 911.6 856.7 6.02% 6791 18.8 1762.9 1800.0 900.4 862.7 4.19% 2554 21.8 1755.4 1800.0 897.1 860.5 4.09% Avg-B 1914.0 1.7 1800.0 3.67% 3070.0 8.7 1800.0 2.87% 1376.0 8.5 1800.0 2.94% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
@@
Table 3: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Extended instances, $L=4$)
@@
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 96 0.0 0.9 1.0 282.4 282.4 0.00% 88 0.1 0.6 0.8 282.4 282.4 0.00% 90 0.3 1.9 2.5 282.4 282.4 0.00% a3-30 106 0.1 11.9 12.2 472.5 472.5 0.00% 101 0.1 9.4 9.7 472.5 472.5 0.00% 101 0.3 8.2 8.7 472.5 472.5 0.00% a3-36 161 0.1 6.9 7.5 510.6 510.6 0.00% 179 0.4 10.6 11.6 510.6 510.6 0.00% 158 0.3 3.6 4.2 510.6 510.6 0.00% a4-48 420 0.3 1798.6 1800.0 661.2 656.5 0.70% 501 1.3 1797.6 1800.0 661.2 657.7 0.52% 404 1.3 1797.6 1800.0 661.2 657.4 0.57% a6-60 859 0.6 1797.8 1800.0 807.1 796.7 1.29% 1623 9.6 1790.0 1800.0 807.1 797.7 1.17% 833 9.8 1788.7 1800.0 807.1 797.6 1.18% a7-70 1502 1.1 1794.6 1800.0 908.5 891.3 1.89% 4266 26.8 1766.3 1800.0 907.9 891.7 1.79% 1341 28.1 1767.3 1800.0 908.5 892.6 1.75% a8-80 1483 1.1 1782.0 1800.0 988.3 958.3 3.03% 5893 66.6 1720.0 1800.0 986.9 958.2 2.90% 1417 62.1 1727.7 1800.0 989.3 959.1 3.06% Avg-A 661.0 0.5 1031.5 0.99% 1807.3 15.0 1031.7 0.91% 620.6 14.6 1030.8 0.94% b3-36 460 0.2 9.0 9.7 530.6 530.6 0.00% 1199 2.5 18.6 22.1 530.6 530.6 0.00% 436 2.2 12.0 14.9 530.6 530.6 0.00% b4-40 910 0.7 733.4 734.9 564.6 564.6 0.00% 2247 18.0 80.7 99.9 564.6 564.6 0.00% 676 18.5 524.5 543.8 564.6 564.6 0.00% b5-50 1686 1.4 1795.4 1800.0 710.8 701.4 1.33% 17112 90.7 1699.1 1800.0 710.8 703.0 1.10% 1420 94.8 1701.4 1800.0 710.8 702.8 1.14% b6-60 3036 3.1 1793.6 1800.0 817.8 799.3 2.26% 22909 177.8 1600.9 1800.0 818.7 799.3 2.38% 2135 185.2 1609.2 1800.0 818.3 804.5 1.69% b7-70 3478 4.1 1778.4 1800.0 837.7 822.7 1.79% 58446 732.6 957.9 1800.0 844.8 821.3 2.78% 2811 713.3 1075.2 1800.0 841.0 824.5 1.97% Avg-B 1914.0 1.9 1228.9 1.08% 20382.6 204.3 1104.4 1.25% 1495.6 202.8 1191.7 0.96% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
Table 4: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Extended instances, $L=6$)
Appendix C Full computational results for MDARP-LPT
Both tables below serve as supplements to Table 7. One point worth noting is that while the instances used in both tables are derived from Cordeau, they are not identical since we modify the number of vehicles and the number of customers (see Table 4 for details).
$S_{RF}$ $S_{FF}$ $S_{PFF}$ F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 45 8.9 0.5 9.4 239.6 239.6 0.00% 40 6.2 0.4 6.6 239.6 239.6 0.00% 35 17.9 0.7 18.7 239.6 239.6 0.00% a3-30 42 7.9 1.4 9.4 357.9 357.9 0.00% 42 9.6 2.1 11.9 357.9 357.9 0.00% 40 10.2 1.2 11.4 357.9 357.9 0.00% a3-36 51 6.5 47.1 53.7 474.8 474.8 0.00% 48 8.0 31.3 39.5 474.8 474.8 0.00% 37 11.1 38.1 49.2 474.8 474.8 0.00% a4-48 182 6.8 217.5 224.7 499.1 499.1 0.00% 166 6.9 179.6 186.8 499.1 499.1 0.00% 124 21.0 131.0 152.2 499.1 499.1 0.00% a6-60 236 7.2 1791.4 1800.0 632.1 624.9 1.14% 210 8.3 1791.2 1800.0 632.1 625.7 1.00% 150 52.3 1557.4 1610.2 632.1 632.1 0.00% a7-70 715 10.3 1788.5 1800.0 646.9 619.6 4.23% 687 13.0 1786.1 1800.0 644.6 622.7 3.40% 364 183.5 1615.6 1800.0 639.9 625.9 2.19% a8-80 613 10.1 1788.7 1800.0 753.6 714.7 5.16% 728 14.2 1784.6 1800.0 747.4 718.0 3.93% 389 192.2 1606.8 1800.0 746.7 720.1 3.56% Avg-A 269.1 8.2 813.9 1.50% 274.4 9.5 806.4 1.19% 162.7 69.7 777.4 0.82% b3-36 203 6.6 148.7 155.5 425.6 425.6 0.00% 227 8.3 479.7 488.2 425.6 425.6 0.00% 133 12.7 181.1 193.8 425.6 425.6 0.00% b4-40 518 9.7 1789.8 1800.0 405.9 402.4 0.85% 502 8.8 156.0 165.1 405.9 405.9 0.00% 260 24.2 132.4 156.8 405.9 405.9 0.00% b5-50 1878 20.9 1777.6 1800.0 497.5 480.1 3.51% 1903 15.6 1783.3 1800.0 494.3 488.4 1.20% 865 44.8 432.3 477.6 494.3 494.3 0.00% b6-60 2537 17.4 1779.9 1800.0 603.1 574.4 4.77% 2334 23.5 1774.6 1800.0 597.8 589.9 1.33% 969 127.6 1230.6 1359.3 596.6 596.5 0.00% b7-70 2310 19.5 1777.3 1800.0 610.8 574.4 5.96% 3287 46.4 1750.2 1800.0 625.7 581.8 7.02% 1145 273.7 1525.2 1800.0 609.0 585.3 3.89% Avg-B 1489.2 14.8 1471.1 3.02% 1650.6 20.5 1210.7 1.91% 674.4 96.6 797.5 0.78% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
@@ Appendix C Full computational results for MDARP-LPT
Table 1: Computational performance of FFF with different fragment sets for MDARP-LPT within 30 minutes (basic instances: $|V|=\lceil n/4\rceil$, $L=4$)

### v3

@@ Appendix B Full computational results of PSFF for DARP-LPT and MDARP-LPT
Tables 13 present the computational performance of PSFF with different fragment sets on the instances from Cordeau [2006]. Specifically, Tables 1 and 2 correspond to DARP-LPT, while Table 3 focuses on MDARP-LPT.
Similarly, Tables 46 evaluate PSFF’s performance on the instances from Riedler & Raidl [2018], with Tables 4 and 5 dedicated to DARP-LPT and Table 6 to MDARP-LPT.
@@
$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time a3-18 302.5 73 0.1 0.0 0.1 64 0.1 0.1 0.2 71 0.1 0.0 0.2 455 0.2 0.0 0.3 1310 0.4 0.0 0.5 a3-30 494.8 93 0.1 0.1 0.2 89 0.2 0.1 0.4 95 0.1 0.1 0.3 1202 0.8 0.1 0.9 2902 0.8 0.1 1.1 a3-36 574.4 121 0.1 0.1 0.3 112 0.1 0.1 0.4 122 0.2 0.2 0.4 1730 1.1 0.1 1.5 4261 1.1 0.2 1.7 a4-48 697.9 318 0.2 0.3 0.7 314 0.9 0.1 1.2 338 0.8 0.1 1.3 4943 5.6 0.2 6.1 14468 5.5 0.4 7.2 a6-60 864.9 601 0.3 0.4 1.1 655 3.4 0.6 4.6 670 1.7 0.4 2.5 14542 12.5 1.5 15.0 48423 17.2 1.9 23.8 a7-70 984.1 1036 0.7 0.8 2.1 1306 4.0 1.9 6.8 1211 2.4 1.0 3.8 19146 22.5 2.1 26.5 57082 21.1 14.4 41.4 a8-80 1058.7 1088 0.6 1.1 2.4 1480 7.1 0.9 8.7 1386 4.5 0.9 5.8 33590 47.3 2.6 52.6 89453 30.3 15.2 56.0 a8-96 1342.2 1465 1.37 2.2 5.0 1684 7.9 4.3 13.3 1709 7.2 6.3 14.3 47210 59.5 5.5 68.8 154853 55.0 20.3 91.6 Avg-A 599.4 0.6 1.5 713.0 1.0 4.5 700.3 1.1 3.6 15352.3 1.5 21.5 46594.0 6.6 27.9 b3-36 576.5 169 0.1 0.1 0.3 166 0.2 0.2 0.5 168 0.2 0.1 0.4 1473 1.4 0.1 1.6 3834 1.2 0.1 1.6 b4-40 635.2 311 0.2 0.1 0.4 212 1.3 0.1 1.6 233 0.5 0.1 0.8 3572 3.9 0.2 4.3 14683 4.8 0.3 6.2 b5-50 789.1 518 0.3 0.7 1.2 804 1.6 0.4 2.5 649 1.4 0.6 2.3 9808 7.4 0.7 8.7 19161 6.3 1.0 9.3 b6-60 927.1 940 0.8 2.3 3.7 1231 2.8 1.6 4.8 1030 1.9 1.1 3.4 12290 15.1 2.1 18.6 28817 9.1 2.1 13.4 b7-70 944.7 960 0.8 4.6 5.9 1678 4.5 3.5 8.7 1188 3.3 2.4 6.0 18758 24.0 2.6 28.5 52404 21.2 3.7 30.0 b8-80 1103.2 1195 0.8 2.3 3.6 1813 5.2 1.5 7.2 1429 4.4 3.0 7.9 26484 43.4 3.0 49.3 71693 27.0 3.4 37.0 b8-96 1294.4 4530 3.1 3.1 7.4 6540 34.9 4.4 40.8 4316 28.5 5.3 34.9 102058 256.3 10.0 273.5 535621 204.4 39.5 291.9 Avg-B 1231.9 1.9 3.2 1777.7 1.7 9.4 1287.6 1.8 7.9 24920.4 2.7 54.9 103744.7 7.2 55.6 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
@@
Table 1: Computational performance of PSFF with different fragment sets for DARP-LPT within 30 minutes (Cordeau’s instances, $L=4$)
@@
$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time a3-18 295.8 73 0.0 0.0 0.1 64 0.5 0.2 1.0 71 0.1 0.0 0.1 470 0.2 0.0 0.3 1310 0.4 0.1 0.6 a3-30 477.0 93 0.1 0.2 0.3 89 0.6 0.9 1.8 95 0.1 0.2 0.7 1250 0.7 0.3 3.1 2902 0.9 0.5 1.8 a3-36 549.8 121 0.1 0.1 0.3 115 0.4 0.1 0.9 122 0.2 0.1 0.4 1951 1.5 0.2 1.9 4261 1.4 0.2 2.3 a4-48 669.4 318 0.2 0.6 1.1 344 1.8 0.5 3.4 339 0.9 0.5 1.6 5818 7.4 1.3 9.5 14468 4.9 4.1 11.3 a6-60 825.0 601 0.3 0.9 1.7 827 5.2 1.1 6.8 673 3.5 0.8 4.7 23568 34.0 4.4 44.4 48423 14.1 19.5 41.3 a7-70 925.7 1036 0.7 6.6 8.3 2157 15.5 3.1 19.5 1233 13.3 2.6 16.5 38512 66.8 10.8 84.7 57082 17.2 126.5 153.1 a8-80 990.8 1088 0.9 8.9 11.1 3028 28.6 7.7 37.3 1390 25.5 6.7 33.0 81310 123.8 24.5 157.7 89453 26.7 317.5 359.6 a8-96 1262.3 1465 1.0 61.7 66.7 2600 25.8 73.2 100.1 1738 26.5 78.2 105.9 80180 128.0 114.4 254.6 154853 49.2 1192.9 1269.1 Avg-A 599.4 9.9 11.2 1153.0 10.8 21.4 707.6 11.1 20.4 29132.4 19.5 69.5 46594.0 207.7 229.9 b3-36 557.9 169 0.1 0.1 0.3 219 1.1 0.2 1.7 172 0.3 0.1 0.5 1723 2.0 0.1 2.3 3857 1.2 0.2 2.1 b4-40 627.4 311 0.2 0.7 1.1 238 5.0 1.0 6.3 244 2.6 1.0 3.8 5033 12.6 1.2 14.5 14683 4.8 1.9 8.8 b5-50 751.9 518 0.4 3.4 4.3 1869 8.6 1.3 10.4 653 5.5 0.9 6.6 31667 48.9 4.6 57.3 19197 5.8 1.6 10.4 b6-60 876.8 940 0.5 3.7 5.0 3818 16.5 2.1 19.3 1079 23.6 2.1 26.2 30029 73.5 10.7 89.8 28817 10.5 5.3 19.7 b7-70 896.4 960 1.0 2.2 4.5 6251 28.7 1.8 31.4 1252 38.3 2.1 41.1 67029 149.1 7.3 163.2 52603 23.0 6.8 38.4 b8-80 1053.8 1195 1.1 12.2 14.5 5896 34.9 14.2 50.1 1481 39.7 10.8 51.4 78389 173.2 14.8 196.8 71693 24.9 14.5 52.4 b8-96 1222.0 4530 3.8 47.1 54.1 51048 902.7 122.7 1029.7 5192 963.1 97.5 1062.4 No 567416 212.9 296.4 597.0 Avg-B 1231.9 9.9 12.0 9905.6 20.5 164.1 1439.0 16.3 170.3 6/7 solved 108323.7 46.7 104.1 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. “No” indicates that no optimal solution is obtained within the 30-minute time limit. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
@@
Table 2: Computational performance of PSFF with different fragment sets for DARP-LPT within 30 minutes (Cordeau’s instances, $L=6$)
@@
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time a3-18 206.7 37 8.1 0.1 8.3 33 7.2 0.1 7.4 35 14.0 0.1 14.3 a3-30 315.7 36 8.3 0.3 8.8 36 7.1 0.3 7.6 36 8.9 0.6 9.8 a3-36 379.9 47 7.4 0.5 10.6 44 16.9 0.6 20.2 47 7.8 0.4 8.7 a4-48 428.7 116 10.0 1.7 13.2 109 11.7 1.8 14.8 119 8.2 1.7 11.0 a6-60 532.2 175 10.2 5.0 17.5 156 10.4 7.4 20.0 176 9.2 5.1 16.5 a7-70 535.4 507 16.3 19.4 40.0 505 17.2 24.3 47.0 528 14.6 21.8 41.2 a8-80 637.3 495 21.2 37.5 65.6 586 20.8 39.5 68.9 576 21.1 42.3 70.4 a8-96 775.6 607 16.0 87.3 116.0 551 29.8 86.5 127.3 602 27.3 81.4 119.9 Avg-A 252.5 19.0 35.0 252.5 20.1 39.1 264.9 19.2 36.5 b3-36 364.1 146 7.5 0.5 8.6 148 8.9 0.5 10.0 147 9.4 0.7 12.1 b4-40 345.1 370 8.4 1.4 11.0 320 9.9 1.3 12.5 322 18.9 1.9 22.0 b5-50 417.0 1649 20.0 20.8 46.7 1682 32.2 14.7 51.4 1457 30.7 18.4 53.7 b6-60 517.9 2269 26.5 30.9 67.3 2084 31.6 21.6 58.8 1788 35.0 23.5 65.2 b7-70 512.4 1541 26.0 62.0 98.6 2049 48.9 27.6 86.4 1493 53.1 48.9 110.6 b8-80 601.0 2062 29.8 116.0 160.7 3133 88.4 96.0 200.7 2267 93.7 101.2 209.1 b8-96 714.8 6520 90.3 345.2 477.6 6582 266.8 170.8 468.5 4485 275.2 304.9 611.1 Avg-B 2079.6 82.4 124.4 2285.4 47.5 126.9 1708.4 71.4 154.8 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
@@
Table 3: Computational performance of PSFF with different fragment sets for MDARP-LPT within 30 minutes (Cordeau’s instances, $L=4$)